OpenStudy (anonymous):

y=3 when x=1/7 write the linear function

OpenStudy (lgbasallote):

proportions?

OpenStudy (anonymous):

y = kx y= 3 and x = 1/7 3 = k * 1/7 k = 7*3 k =21 y = 21x

OpenStudy (lgbasallote):

how do you know it's not y = k/x?

OpenStudy (anonymous):

Linear